OpenStudy (anonymous):

solution set for quadratic equation (x+9)^2=-25

OpenStudy (anonymous):

Distributing out, it becomes x^2+18x+106 = 0

OpenStudy (anonymous):

Using the quadratic formula, we obtain: x= -9 (+/-) 5 i, where i is sqrt(1)

OpenStudy (anonymous):

The roots are complex.

myininaya (myininaya):

take square root of both sides x+9=plus or minus sqrt{-25} x=-9 plus or minus sqrt{-25} x=-9 plus or minus 5i x=-9+5i or x=-9-5i